3-minute song is sampled at 44.1 KHz with 65,536 quantization levels in mono (one

channel). How big would the mp3 be given an mp3 is 1/11 of the uncompressed size?
Answer #1

3 minutes = 3 * 60 seconds = 180 seconds

Sampling rate = 44.1 KHz = 44100 Hz

Quantization levels = 65,536 = 216

So, bits required to represent 65,536 quantization levels = 16

Number of channel = 1 (mono)

File size = ( sampling rate * bits per sample * channel number * duration ) bits

= (44100 * 16 * 1 * 180) bits = 127008000 bits = 15876000 Bytes = 15503.906 KB = 15.14 MB

So, the uncompressed size of the mp3 file is 15.14 Megabytes (approximately).

With 11:1 compression ratio the mp3 file size will be = (15.14/11) MB = 1.376 MB

So, the compressed size of the mp3 file is 1.376 Megabytes (approximately).

Hope this helps.